OpenStudy (anonymous):

7. Evaluate 1/2a^–4b^2 for a = –2 and b = 4. A. 2 B. 1/2 C. 3 D. 4

OpenStudy (anonymous):

Is it \( \huge \frac 12a^{–4}b^2 \)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then the answer is B. 1/2.

OpenStudy (anonymous):

Convert everything to power of \( 2 \) and then use laws of indices.

OpenStudy (ash2326):

we have \[1/2 a^{-4} b^2\] substitute a=-2 and b=4 we get \[1/2 (-2)^{-4} (4)^2\] or \[\frac{16}{2*16}\] cancelling the 16 from numerator and denominator \[\frac{1}{2}\]